OpenStudy (anonymous):

a bag contains 4 green marbles, 6 red marbles, 14 orange marbles, 5 brown marbles, and 8 blue marbles. you chose a marble, replace it, and choose again. what is P(red, then blue)? 48/1369 14/1369 27/37 14/37

OpenStudy (warriorz13):

So I would first rule out A,B because of the 1369 when you don't even have a # over 14

OpenStudy (anonymous):

so 14/1369

OpenStudy (warriorz13):

But what is YOUR guess

OpenStudy (anonymous):

that was my guess

OpenStudy (warriorz13):

No, I would rule those out because 4+6+14+5+8 is only 37 but my guess is D

OpenStudy (warriorz13):

But wait a sec @IsaiahCC100 @jojo9999 @ganeshie8 @texaschic101

OpenStudy (anonymous):

NO this is probability much different than just adding chances are much much lower than 2/37

OpenStudy (warriorz13):

?????????????????????

OpenStudy (anonymous):

its either 48/1369 or 14/1369

OpenStudy (warriorz13):

But there arte no choices lower then 14/37

OpenStudy (anonymous):

48/1369 and 14/1369 are much much lower than it

OpenStudy (anonymous):

IT WOULD BE THOSE WARRIORZ

OpenStudy (texaschic101):

4 + 6 + 14 + 5 + 8 = 37 chance of drawing a red are 6/37 chances of drawing are 8/37 6/37 x 8/37 = 48/1369 I am not the best with probabilities, so I do not even know if this is correct...this is just a guess

OpenStudy (anonymous):

Thank you texas

OpenStudy (texaschic101):

oops..chance of drawing blue are 8/37

OpenStudy (anonymous):

People think probability is easy when its one of the hardest classes ever.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Boom

OpenStudy (warriorz13):

But it is easy LMFAO I passed that section of math with a 95%

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

@Warriorz13 If its so easy how did you not do it right? I mean its easy man!

OpenStudy (warriorz13):

But you just said it'shard-___-

OpenStudy (anonymous):

But you said its easy

OpenStudy (texaschic101):

it can be confusing...because sometimes you add them and sometimes you multiply them

OpenStudy (anonymous):

Exactly but the actual class isnt just confusing its mind blowing

OpenStudy (texaschic101):

my mind got blown..oh...that just sounds bad

OpenStudy (anonymous):

Only in your mind.. only in yours.

OpenStudy (texaschic101):

lol